Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to sell one type of carburetor to Troy Engines, Ltd., for a cost of $39 per unit. To evaluate this offer, Troy Engines, Ltd., has gathered the following information relating to its own cost of producing the carburetor internally:

Per Unit 21,000 Units
per Year
Direct materials $ 18 $ 378,000
Direct labor 11 231,000
Variable manufacturing overhead 3 63,000
Fixed manufacturing overhead, traceable 3 * 63,000
Fixed manufacturing overhead, allocated 6 126,000
Total cost $ 41 $ 861,000

*One-third supervisory salaries; two-thirds depreciation of special equipment (no resale value).

Required:

1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 21,000 carburetors from the outside supplier?

2. Should the outside supplier’s offer be accepted?

3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $210,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 21,000 carburetors from the outside supplier?

4. Given the new assumption in requirement 3, should the outside supplier’s offer be accepted?
Answer #1
1
Per unit Total 21000 units
Make Buy Make Buy
Direct materials 18 378000
Direct labor 11 231000
Variable manufacturing overhead 3 63000
Fixed manufacturing overhead traceable 1 21000
Purchase cost 39 819000
Total 693000 819000
Difference in favor of making = 693000-819000 = $126000
Financial (disadvantage) (126000)
2
NO, Reject the offer
3
Make Buy
Total cost 693000 819000
Opportunity cost 210000
Total relevant cost 903000 819000
Difference in favor of buying = 903000-819000 = $84000
Financial advantage 84000
4
Yes, Accept the offer